Find the equation of a line that is parallel to y = -x/2 + 4 and that passes through the point (4, 0).
Answer by stanbon(75887)

Find the equation of a line that is parallel to y = -x/2 + 4 and that passes through the point (4, 0).
-----
The given line has slope = -1/2.
----
Every line parallel to it must have slope = -1/2
-----------------------
Form: y = mx + b
Solve for "b":
0 = (-1/2)4 + b
b = 2
------
Equation:
y = (-1/2)x + 2
========================
Cheers,
Stan H.
===================